OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. x + 2y = 12 -x = -y - 6

OpenStudy (anonymous):

What, another one...?

OpenStudy (anonymous):

just tell me in which u r having trouble?

OpenStudy (anonymous):

It doesnot makes sense if we go on solving all ur problems coz u will learn nothing

OpenStudy (anonymous):

Let us help u

OpenStudy (anonymous):

it's not for me. It's for my brother.

OpenStudy (anonymous):

Ahh I can see that, well then x=y+6 Now substitute into 1 y+6+2y=12 3y=6 y=2 x=2+6=8

OpenStudy (anonymous):

(y+6)+2y = 12 y+6+2y = 12 3y = 12-6 y = 6/3 y = 2 substitute y in equation x+2(2) = 12 x+4 = 12 x = 12-4 x = 8 so, x=8,y=2

OpenStudy (amistre64):

lol ... tell your brother to do his own work ;)

OpenStudy (anonymous):

it will be like this we will multiply both side of this : -x=-y+16 into minus so we will get x=y+6 Now it is time to sustitute in the question: x+2(2)=12 x+4=12 x=8 then we will get x=8,y=2

OpenStudy (amistre64):

id just use it as is: x + 2y = 12 -x = -y - 6 --------------- 2y = -y +6 ; +y 3y = 6 ; /3 y = 2 ; use that to determine the value of x